OpenStudy (anonymous):

int (x^2)(cosx)dx please help :)

OpenStudy (xapproachesinfinity):

so where did you stuck

OpenStudy (anonymous):

hmmm ok ive got u=cosx du=-sinx dv=x^2dx v=x^3/3

OpenStudy (anonymous):

\[x ^{3}cosx-\int\limits_{}^{}-x^3sinxdx\]

OpenStudy (anonymous):

well x^3sinxdx/3

OpenStudy (xapproachesinfinity):

eh you are complicating things up why don't you pick u=x^2 and dv=cosx

OpenStudy (anonymous):

ah yeah that is a bit cleaner

OpenStudy (xapproachesinfinity):

we are trying to avoid x^3 and sort of expressions

OpenStudy (anonymous):

so v=sinx right

OpenStudy (xapproachesinfinity):

now redo the Integral by part

OpenStudy (xapproachesinfinity):

no! if dv=cosx v=-sinx

OpenStudy (anonymous):

i need to take a second substitution... so im guessing u-2x and dv=sinx?

OpenStudy (xapproachesinfinity):

oh hold on sorry you are right

OpenStudy (anonymous):

i always get confused with the sinx and cosx derivatives and which is negative... lol

OpenStudy (xapproachesinfinity):

yes! go for it now

OpenStudy (anonymous):

ok so du=xdx and v=-cosx

OpenStudy (xapproachesinfinity):

Don't worry we all do such mistakes just recheck your stuff by differentiating and see if that's true

OpenStudy (anonymous):

do i need to do a third sub here?

OpenStudy (anonymous):

i end up with x^2sinx+2xcosx-int(-xcosxdx)

OpenStudy (xapproachesinfinity):

v=sinx not cosx i said you are right

OpenStudy (anonymous):

wait i meant v=-cosx for the second substitution

OpenStudy (xapproachesinfinity):

so u=x^2 dv=cosx du=2x v=sinx just make sure this is what you have ok

OpenStudy (anonymous):

for the first sub yes

OpenStudy (anonymous):

then with that i get x^2sinx-intsinx2xdx

OpenStudy (xapproachesinfinity):

oh second int by part you -2"intxsinx"

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ah do i have to pull the 2 out or can i just make u=2x

OpenStudy (anonymous):

nah probably best if i take the 2 out huh

OpenStudy (xapproachesinfinity):

pull it out! make u=x and dv=sinx

OpenStudy (xapproachesinfinity):

yes! make thing easier! alway s think about an easier method

OpenStudy (anonymous):

so v=-cosx

OpenStudy (xapproachesinfinity):

correct!

OpenStudy (anonymous):

ok so

OpenStudy (xapproachesinfinity):

now you will see that it is done already just make sure you add all what we did together

OpenStudy (anonymous):

\[x^2sinx-2[-xcosx-\int\limits_{}^{}-cosxdx]\]

OpenStudy (anonymous):

is that right what i did with the 2? put everything after it in brackets?

OpenStudy (xapproachesinfinity):

Correctt -2 is a factor to all of that second part

OpenStudy (anonymous):

ok so final answer

OpenStudy (anonymous):

\[x^2sinx+2xcosx+2sinx+C\]

OpenStudy (xapproachesinfinity):

you missed a negative sign should be -2sinx

OpenStudy (anonymous):

oh yeah i see it now thank you

OpenStudy (xapproachesinfinity):

everything else is good! and you are welcome

OpenStudy (xapproachesinfinity):

in integration by part pay real attention to details you might forgot things so when you clean up the miss return to your work and see if that's true! comprend!

OpenStudy (xapproachesinfinity):

i checked by differentiating it and all is clean and good! YaY

OpenStudy (xapproachesinfinity):

mess* (correction)